OpenStudy (anonymous):

A child lies on the ground and looks up at the top of a 9-ft tree nearby. The child is 13 ft away from the tree. What is the angle of elevation from the child to the top of the tree? Round to the nearest whole degree.

OpenStudy (anonymous):



OpenStudy (anonymous):

here tan theta=9/13

OpenStudy (anonymous):

35 degree

jimthompson5910 (jim_thompson5910):

tan(theta) = opp/adj tan(theta) = 9/13 theta = arctan(9/13) theta = 34.69515 theta = 35 degrees (round to the nearest whole number)

jimthompson5910 (jim_thompson5910):

You can find the "arctan" button by hitting "2nd" then the tangent button (if you have a TI 83 or TI 84 calculator)

OpenStudy (anonymous):

ha ha i googled it!!:D